OpenStudy (anonymous):

What is the domain and range of th function y=x^(1/3)?

OpenStudy (gabylovesyou):

https://www.purplemath.com/modules/fcns2.htm

OpenStudy (anonymous):

The domain is given by the values of X in which the expresion y = x(1/3) makes sense. Know that \[x ^{(1/3} = \sqrt[3]x{}\] And observe that you can got the 3sqrt of any number. No matters if it is positive or negative. Then, the domain will be all the numbers in R (-infinity, + infinity)